OpenStudy (anonymous):

Karli's shadow is 60 inches long. A nearby fire hydrant casts a shadow 40 inches long. If Karli is 48 inches tall, what is the height of the fire hydrant?

rishavraj (rishavraj):

use unitary method

OpenStudy (anonymous):

the unitary method?? whats that?

OpenStudy (anonymous):

wait would it be 28 inches?

rishavraj (rishavraj):

no answerbis somthng else

OpenStudy (anonymous):

so how do I find the answer? what are the steps?

rishavraj (rishavraj):

see when shadow is of 60 inches the body is of 48 inch... (think opposite).. then its clear that when body will be of 48/60 inches then the lenght of shadow would be 1 inch

OpenStudy (anonymous):

hmmm, that's really confusing lol

rishavraj (rishavraj):

have u studied solving via unitary method

OpenStudy (anonymous):

no im in this credit recovery program where they just throw the problems at u and then u just figure em out haha, so I don't know any of this really

rishavraj (rishavraj):

visit the website http://www.algebraden.com/unitary_method.htm

rishavraj (rishavraj):

have u understood that object of lenght 48/60 inches will cast shadow of 1 inch

OpenStudy (anonymous):

not really to be honest

rishavraj (rishavraj):

see solve a question If a bus can carry 28 passengers..then how many buses will be required to carry 224 passengers??

OpenStudy (anonymous):

8

rishavraj (rishavraj):

yeah now solve this question the same way... forget that there a hydrant or karli

OpenStudy (anonymous):

oh so u get.8 then just round up?

rishavraj (rishavraj):

i didnt get it

OpenStudy (anonymous):

I divided 48/60 and got .8

rishavraj (rishavraj):

u forgot somthinng to multiply I guess

rishavraj (rishavraj):

just have a look over the solution of the question which i asked u

OpenStudy (anonymous):

find the ratio that if you multiply the shadow, it gives you the height. known ratio = 48/60 and you got .8 multiply 0.8 by karls shadow 0.8 * 60 = 48 you get his height, so we know the raito works. so now you have the ratio you can multiply this by any shadow and get the height. height = ratio * shadow height = 0.8 * 40